what is the solution to 7/2x+7/3y=42 , 1/4x+1/3y=4
Answer by richwmiller(17219)

Are x and y part of the numerators or denominators?
If part of the denominator x=1/8 and y=1/6
if part of the numerator x=8 and y=6